OpenStudy (anonymous):

how to factorize -(a^2 +b^2)? does it equal to -(a^2+2ab+b^2)?

OpenStudy (anonymous):

Ever heard of the "difference of two squares" rule? It states that: \[\Large {a^2} - {b^2} = (a + b)(a - b)\] In your case: \[\Large \begin{array}{l} - ({a^2} + {b^2})\\ = {b^2} - {a^2}\\ = ... \end{array}\]

OpenStudy (anonymous):

thx=]

OpenStudy (anonymous):

No worries =) A medal would be nice though :-P